OpenStudy (anonymous):

Two cyclist A and B traveled the same distance of 120 kilometers. A traveled 4 kilometers per hour faster than B and covered the distance in one hour less than B. Find the rate of each.

OpenStudy (anonymous):

(x+4)\[\frac{ 120 }{ x+4 }=\frac{ 120 }{ x }+1\] x is the rate at which B is travelling (ie. xkm/h)

OpenStudy (anonymous):

There SOLVE FOR x

OpenStudy (anonymous):

Wait it's incorrect for some reason. I made a mistake.

OpenStudy (anonymous):

Tried solving it, but nothing came out.

OpenStudy (anonymous):

120/(x+4) =120/x +1 120/(x+4) =(120+x)/x 120x =(120+x)(x+4) 120x =120x +480 +x^2 +4x x^2 +4x +480 =0

OpenStudy (anonymous):

That's the problem.

OpenStudy (anonymous):

which has NO real SOLUTION

OpenStudy (anonymous):

It is should be 120/(x+4) =120/x -1

OpenStudy (anonymous):

Beacuse A takes 1 hour less than B

OpenStudy (anonymous):



OpenStudy (anonymous):

U need to solve it

OpenStudy (anonymous):

Yeah. A bit of confusion but we got there.

OpenStudy (anonymous):

x is a imaginary number.

OpenStudy (anonymous):

No. Read suarav's post.

OpenStudy (anonymous):

it's not +1 it's -1

OpenStudy (anonymous):

-2 + 21.82i or -2-21.82i

OpenStudy (anonymous):

\[\frac{ 120 }{ x+4 }=\frac{ 120 }{ x }-1\]

OpenStudy (anonymous):

I previously wrote +1

OpenStudy (anonymous):

instead of -1

OpenStudy (anonymous):

oh really sorry, I have a mistake. x=24.

OpenStudy (anonymous):

That's incorrect still. Try again.

OpenStudy (anonymous):

really?? . so, x=20, right??? Don't tell me it's still incorect ..:(

OpenStudy (anonymous):

Correct. Well done.